1.Find the general solution to the following ODE's
a). y'' +y= sec^2t
b). x^2y'' +3xy'+3y=0
Answer #1

I can y+y = secat L (D2+12y = sec?t A. E D2+1=0 => D= . . Cif a cost + G sint U= cost, vs Sint vi , cost sint la cos+ sin?t=